# Projection effects in barred galaxies causing wrong interpretation of radial flows

### v1

@@
Galaxy disks in rotation are sometimes the site of radial flows, especially in their gas component. It is important to estimate the outflows, due to AGN or supernovae feedback, or inflows due to bar gravity torques. However, these radial flows may be confused with non-circular motions, which are quite frequent in the center of galaxy disks. We use a simulated giant, barred spiral galaxy from the GalMer database to study the non-circular motions induced by the bar. Our goal is to identify the spurious radial flows that kinematics modeling algorithms can detect, assuming circular orbits for the gas. Using mock data of a strongly barred galaxy, we quantify the radial velocities computed by the 3D-Barolo algorithm for different disk inclinations and several bar orientations in the plane of the sky: along the major and minor kinematic axes and at $\pm 45^{\circ}$ from them. Our results show that projection effects cause kinematics modeling algorithms to confuse the radial component of velocity due to elliptical orbits with significant radial flows with mean values up to 84 km s-1  within the bar region. The computed rotation curve is also wrongly estimated inside the bar region, by as much as 150 km s-1for the highest inclination.
@@ 1 Introduction
Bars are associated with an $m=2$ instabilities with two-arm spiral waves, producing a bisymmetric gravitational potential (Buta & Combes, 1996). They represent a long-lived, quasi-stationary density wave arising from the superposition of leading waves, which transfer angular momentum inward, and trailing waves, which transfer angular momentum outward. Bars exchange angular momentum with the outer disk and dark matter halo through resonant interactions, particularly at the corotation and outer Lindblad resonances (OLR), which allows them to grow and slow down. Bars induce spiral structure in the gas component, which is dissipative: gas cloud collisions prevent the gas to follow crossing orbits like stars (Sanders & Huntley, 1976; Buta & Combes, 1996). Bars and spirals are then out of phase, and bars exert gravity torques on the gas. Inside corotation, these torques are negative, and the gas loses angular momentum to the benefit of the bar. This leads to gas accretion of central supermassive black holes (SMBHs). In fact, in active galactic nuclei (AGNs), the gas angular momentum has to be reduced by 5 orders of magnitude when infalling from a typical galactocentric radius of 3 kpc to the last stable orbit, over a timescale much shorter than that of the known efficient mechanisms (Combes, 2023).
@@
In specific regions of barred galaxies, the gas can be trapped in rings at resonances, when there is a rational multiple between motion frequencies of the particles and the bar wave, in the bar rotating frame (see Contopoulos & Grosbol (1989) for a review on stellar orbits in barred galaxies). When the orbital frequencies satisfy the condition for the $m=2$ Lindblad resonances (for a bar wave), i.e. when $\Omega\pm\kappa/2=\Omega_{b}$ where $\Omega$ is the circular frequency, $\kappa$ the epicyclic frequency, and $\Omega_{b}$ the bar pattern speed, the orbits close after two radial oscillations (Buta & Combes, 1996) in the bar rotating frame. Inside corotation, the ILR, associated with a negative sign in the previous equation, plays a role in damping waves that might prevent bar formation, or weaken the bar. Inside the corotation region (CR), stars follow $x_{1}$ orbits: periodic elongated orbits that are parallel to the bar (Athanassoula, 1992). These orbits support the bar structure. Stellar orbits orientation alternates between parallel and perpendicular ($x_{2}$ orbits) to the bar at each resonance (Contopoulos & Papayannopoulos, 1980; Athanassoula, 1992). At resonance radii, gas follows stellar orbits so it also has elongated orbits at the ILR in barred galaxies. At other radii, the inclination of the gas orbits varies, they do not change drastically, hence the spirals (Combes et al., 2014).
Combes et al. (2014) were the first to observe a trailing spiral structure driving gas inward within the ILR ring of a nuclear bar in NGC 1566, showing what kind of morphological features allow the gas to fuel the central SMBH. In fact, negative torques have been computed in the central 300 pc, removing angular momentum from the gas under the SMBH’s influence. At a resonance, the sign of the torque exerted by the bar on the gas changes, causing gas to accumulate in rings. Inside the ILR, the central mass concentration can modify the potential, changing the precession rate and reversing the torque sign again, so only trailing spirals persist, funneling gas inward toward the nucleus (Combes, 2023). This also weakens the bar (Bournaud & Combes, 2002).
@@
Observational programs such as NUclei of GAlaxies (NUGA) and the Galaxy Activity, Torus, and Outflow Survey (GATOS) investigate the gas dynamics and activity in the close environment of AGN, at the sub-kpc scale. While NUGA focused on circumnuclear molecular gas and gravitational torques, GATOS targets ionized and molecular outflows on similar scales, probing the interplay between inflow and feedback. The computation of gravity torques from the bar have revealed negative torques inside CR, leading to gas inflow toward the nucleus (Casasola et al., 2011). These systems often display embedded structures, such as ’bars-within-bars’ configurations (Casasola et al., 2008). In several galaxies from the GATOS survey, ionized gas outflows have been observed and characterized (Davies et al., 2024; Zhang et al., 2024; Hermosa Muñoz et al., 2024). Ionized outflows have a complex geometry and not necessarily match the molecular outflow (Davies et al., 2024). They could be triggered by radio jets (Zhang et al., 2024) and cause positive feedback inducing star formation in the interstellar medium of the galaxy (Hermosa Muñoz et al., 2024). In order to properly quantify these radial flows, it is important to distinguish and separate in the gas kinematics what is due to non-circular motions, from the actual flows. Randriamampandry et al. (2016) have shown with simulated mock galaxies, analysed with the algorithms ROTCUR and DiskFit, that the gas kinematics can be strongly biased, and a true rotation curve difficult to obtain.
@@ 2 Data
We based our study on a gas-rich giant spiral barred galaxy (gSb) from the GalMer database. This database has been built using smoothed-particle hydrodynamics (SPH) simulations to study the impact of intergalactic interactions on galaxy evolution, but it contains isolated galaxy runs as control samples. For the purpose of our analysis we selected one of these isolated galaxy runs. We chose one with 160 000 particles (run #1030). We wanted well-defined spiral arms and a strong bar so as in Randriamampandry et al. (2016) we analyzed the snapshot at $T=500$ Myrs. From now on, the model will be referred to as gSb. The properties of gSb are given in Table 1.
@@
$M_{\rm{B}}\ [10^{9}M_{\odot}]$ 11.5
$M_{\rm{H}}\ [10^{9}M_{\odot}]$ 172.5
$M_{*}\ [10^{9}M_{\odot}]$ 46
$M_{\rm{g}}\ [10^{9}M_{\odot}]$ 9.2
@@
Table 1: Fundamental parameters for gSb: mass of the different components, scale lengths and number of gas particles from Chilingarian et al. (2010).
@@
From the GalMer website [^1] http://galmer.obspm.fr, we obtained masses, positions and velocities of gas particles. The galaxy is symmetrized by adding the same particles at opposite positions (x,y,z)(-x,-y,-z) and opposite velocities $(-v_{x},-v_{y},-v_{z})$. Then, spectral cubes are created only with the gas particles with each pixel measuring 200 pc $\times$ 200 pc and using a spectral resolution of 12 km s-1. Initially, the galactic disk is in the xyx-y plane and the line-of-sight (LoS) is in the $z$ direction. We create one cube every 10 for an inclination angle $i_{\rm{true}}\in[30^{\circ},80^{\circ}]$, the interval recommended by Di Teodoro & Peek (2021). In order to study symmetrical configurations with respect to the LoS, we vary the inclination orientation which is equivalent to changing the nearest/farthest side. To reproduce observations we smooth the obtained cube using a Gaussian beam with a full width at half maximum (FWHM) of 0.3 kpc (corresponding to $\arcsec=1$ kpc for a galaxy at a distance of $D=206.3$ Mpc). This FWHM corresponds to a maximum baseline of $B_{\rm max}=724$ m at $\nu=350$ GHz, consistent with ALMA CO(3–2) observations in a compact configuration. As an example, this typical emission line was used by Audibert et al. (2019) to observe a trailing spiral inside the ILR ring of the bar in NGC 613.
@@ 4 Kinematics modeling with the bar along the kinematic axes
This allows us to evaluate how these limiting configurations affect 3D-Barolo’s recovery of the geometrical and kinematic parameters plotted in Fig 2.
Figure 2: Each column corresponds to one bar orientation of Fig. 1. Left column: bar along the minor axis and right column: bar along the major axis. Parameters fitted by 3D-Barolo with a free PA and a free inclination angle. Each marker correspond to one ’true’ inclination value $i_{\rm{true}}\in\pm[30^{\circ},80^{\circ}]$, the one we input to create the cube. The different parameters are plotted as function of the galactocentric radius. First row: the PA obtained at the end of step 1. Second row: the inclination angles obtained at the end of step 1. Third row: rotation curves obtained after step 2. The black dashed curve represents the expected circular velocity (Eq. 6) computed from the gravitational potential (Eq. 4). Fourth row: Radial velocity with the sign convention: negative towards the center.
@@
We fit both of the geometrical parameters: the PA $\phi$ and the inclination angle $i$ and the kinematic parameters: the tangential velocity $V_{\rm rot}$, the velocity dispersion $V_{\rm disp}$ and the radial velocity $V_{\rm rad}$, plotted in Fig 2.
@@
Because radial motions project onto the minor axis, the bar–driven radial component is absent in the recovered kinematics when the bar is aligned with the major axis (right column, last row, panel of Fig. 2), whereas a non-zero radial component appears when the bar is along the minor axis (left column, last row, panel of Fig. 2). These amplitudes remain lower than in the $\pm 45^{\circ}$ configurations (last row of Fig. 47), explaining our choice of dedicated analysis of those orientations in Sect. 5.
Finally, when radial contribution is almost zero, that is when the bar is along the major axis, 3D-Barolo cannot explain the bar effects by fitting radial velocities, so it shifts the PA by up to 55 within the bar region.
@@ 5 Kinematics modeling with the bar at $\pm 45^{\circ}$ from the kinematic axes
  • a free PA and a fixed inclination angle (Sect 5.3).
@@ 5.1 Results with a free PA and a free inclination angle
Figure 4: Same as Fig. 2 but each column corresponds to one configuration presented in the rows of Fig 3.
@@
The lower the inclination, the further the PA estimated by 3D-Barolo in the region $R<4$ kpc is from the expected PA of $\phi=270^{\circ}$ (first row of Fig. 4). Outside this region, the computed PA corresponds to the correct value we used as a fixed input in Sect. 5.2, $\phi=270^{\circ}$. However, inside the bar, 3D-Barolo tends to oppose the PA to the bar orientation when it is at $\pm 45^{\circ}$. This is likely due to the influence of the spiral structure around the bar. For a highly inclined disk ($i_{\rm true}=80^{\circ}$ in all configurations and also $i_{\rm true}=70^{\circ}$ in configurations (c) and (d)), 3D-Barolo computes the right PA even in the bar region. The bar is less observable in almost edge-on disks and the algorithm seems to ignore its effect on the geometrical parameters in this case.
@@ 5.3 Results with a free PA and a fixed inclination angle
Fixing the inclination produces a stronger overestimation of the tangential velocity at high inclination in configurations (c) and (d) (second row of Fig. 7). Once the inclination can no longer adjust, the fit compensates for the bar-induced non-circular motions by artificially increasing $V_{\rm rot}$. In the other cases, the rotational curves are similar than the ones obtained with a free $i$ (third row of Fig 4).
@@ 6 Discussion
The first limitation of our study concerns the algorithm itself. As noted by Di Teodoro & Peek (2021), 3D-Barolo assumes an axisymmetric disk, which is not an adequate description for a barred spiral galaxy. When used on observations, 3D-Barolo usually fits the geometrical parameters of the galactic disk but even in the best cases, where the PA or the inclination angle are known and fixed to a constant value, one cannot escape the bar’s effects.
@@
In most cases, when a Bézier curve fits the inclination angle, the mean value computed by 3D-Barolo outside the bar region is closer to the true value than when the PA is fixed at $\phi=270^{\circ}$ (first row of Fig 6) than when it is left free (second row of Fig. 4). However, there are some exceptions in configurations (c) and (d). 3D-Barolo tends to underestimate the median inclination more compared to the values obtained when the PA is free to vary (between 1.9 and 5.1 versus 1.1 and 4.3).
The rotation curves are similar whether the PA is left free (third row of Fig. 4) or fixed (the second row of Fig 6) outside $R>4$ kpc. At $i_{\rm true}=80^{\circ}$, there are unexpected high values of $V_{\rm rot}$ in configurations (b) and (d) (second row of Fig. 6) around $R\sim 3$ kpc that are not there when the PA is free (third row of Fig. 4) as the computed inclination was a constant median value.
The apparent radial velocities obtained by 3D-Barolo in the bar region when the PA is left free (fourth row of Fig. 4), are lower than with a fixed PA of $\phi=270^{\circ}$ (third row of Fig 6).
@@
For highly inclined disk, the apparent radial velocities decrease, reaching values close to 0 for $i_{\rm{true}}=80^{\circ}$ in certain cases. The values change a lot with the configurations and the input parameters combination. This confirms that the bar affects the recovery of both geometrical and kinematic parameters, introducing substantial uncertainty in the apparent radial velocities inferred by 3D-Barolo. The low radial velocities force the model to absorb the bar-driven non-circular motions into the tangential component, producing the artificially high values of $V_{\rm{rot}}$ at this inclination (second rows of Fig. 6 and Fig. 7). At this high inclination, 3D-Barolo attempts to fit circular orbits, effectively ignoring the non-circular motions produced by the bar.
@@
We showed that projection effects can induce a misinterpretation of radial motions in barred spiral galaxies. Identifying non-axisymmetric structures is therefore a prerequisite before interpreting apparent radial motions as genuine inflows or outflows. In particular, true radial inflows of gas are expected to be weak on average and can only be reliably measured through gravitational torque computations.

### v2

@@
Galaxy disks in rotation are sometimes the site of radial flows, especially in their gas component. It is important to estimate the outflows, due to AGN or supernovae feedback, or inflows due to bar gravity torques. However, these radial flows may be confused with non-circular motions, which are quite frequent in the center of galaxy disks. We use a simulated giant, barred spiral galaxy from the GalMer database to study the non-circular motions induced by the bar. Our goal is to identify the spurious radial flows that kinematics modeling algorithms can detect, assuming circular orbits for the gas. Using mock data of a strongly barred galaxy, we quantify the radial velocities computed by the 3D-Barolo algorithm for different disk inclinations and several bar orientations in the plane of the sky: along the major and minor kinematic axes and at $\pm 45^{\circ}$ from them. Our results show that projection effects cause kinematics modeling algorithms to confuse the radial component of velocity due to elliptical orbits with significant radial flows with mean values up to 92 km s-1  within the bar region. The computed rotation curve is also wrongly estimated inside the bar region, by as much as 150 km s-1for the highest inclination.
@@ 1 Introduction
Bars are associated with an $m=2$ instabilities with two-arm spiral waves, producing a bi-symmetric gravitational potential (Buta & Combes, 1996). They represent a long-lived, quasi-stationary density wave arising from the superposition of leading waves, which transfer angular momentum inward, and trailing waves, which transfer angular momentum outward. Bars exchange angular momentum with the outer disk and dark matter halo through resonant interactions, particularly at the corotation and outer Lindblad resonances (OLR), which allows them to grow and slow down. Bars induce spiral structure in the gas component, which is dissipative: gas cloud collisions prevent the gas to follow crossing orbits like stars (Sanders & Huntley, 1976; Buta & Combes, 1996). Bars and spirals are then out of phase, and bars exert gravity torques on the gas. Inside the corotation region (CR), these torques are negative on gas weighted average, and the gas loses angular momentum to the benefit of the bar. In fact, in active galactic nuclei (AGNs), the gas angular momentum has to be reduced by 5 orders of magnitude when infalling from a typical galactocentric radius of 3 kpc to the last stable orbit, over a timescale much shorter than that of the known efficient mechanisms leading to gas accretion onto the central supermassive black hole (SMBH) (Combes, 2023). Bars also exchange angular momentum with the outer disk and the dark matter halo through resonant interactions, particularly at the corotation radius and OLRs, which allows them to grow and slow down (Bournaud & Combes, 2002).
@@
In specific regions of barred galaxies, the gas can be trapped in rings at resonances, when there is a rational multiple between motion frequencies of the particles and the bar wave, in the bar rotating frame (see Contopoulos & Grosbol (1989) for a review on stellar orbits in barred galaxies). When the orbital frequencies satisfy the condition for the $m=2$ Lindblad resonances (for a bar wave), i.e. when $\Omega\pm\kappa/2=\Omega_{b}$ where $\Omega$ is the circular frequency, $\kappa$ the epicyclic frequency, and $\Omega_{b}$ the bar pattern speed, the orbits close after two radial oscillations (Buta & Combes, 1996) in the bar rotating frame. Inside corotation, the ILR, associated with a negative sign in the previous equation, plays a role in damping waves that might prevent bar formation, or weaken the bar. Stellar orbits orientation alternates between parallel ($x_{1}$ orbits) and perpendicular ($x_{2}$ orbits) to the bar at each resonance (Contopoulos & Papayannopoulos, 1980; Athanassoula, 1992). These orbits support the bar structure. At other radii than the resonance ones, the inclination of the gas orbits varies, they do not change drastically, hence the spirals (Combes et al., 2014).
Combes et al. (2014) were the first to observe a trailing spiral structure driving gas inward within the ILR ring of a nuclear bar in NGC 1566, showing what kind of morphological features allow the gas to fuel the central SMBH. In fact, negative azimuthally-averaged and gas-weighted torques have been computed in the central 300 pc, removing angular momentum from the gas under the SMBH’s influence. At a resonance, the sign of the torque exerted by the bar on the gas changes, causing gas to accumulate in rings. Inside the ILR, the central mass concentration can modify the potential, changing the precession rate and reversing the torque sign again, so only trailing spirals persist, funneling gas inward toward the nucleus (Combes, 2023). This also weakens the bar (Bournaud & Combes, 2002).
@@
Observational programs such as NUclei of GAlaxies (NUGA) and the Galaxy Activity, Torus, and Outflow Survey (GATOS) investigate the gas dynamics and activity in the close environment of AGN, at the sub-kpc scale. While NUGA focused on circumnuclear molecular gas and gravitational torques, GATOS targets ionized and molecular outflows on similar scales, probing the interplay between inflow and feedback. The computation of gravity torques from the bar have revealed negative gas-weighted averaged torques inside CR, leading to gas inflow toward the nucleus (Casasola et al., 2011). These systems often display embedded structures, such as ’bars-within-bars’ configurations (Casasola et al., 2008). In several galaxies from the GATOS survey, ionized gas outflows have been observed and characterized (Davies et al., 2024; Zhang et al., 2024; Hermosa Muñoz et al., 2024). Ionized outflows have a complex geometry and not necessarily match the molecular outflow (Davies et al., 2024). They could be triggered by radio jets (Zhang et al., 2024) and cause positive feedback inducing star formation in the interstellar medium of the galaxy (Hermosa Muñoz et al., 2024). In order to properly quantify these radial flows, it is important to distinguish and separate in the gas kinematics what is due to non-circular motions, from the actual flows. Randriamampandry et al. (2016) have shown with simulated mock galaxies, analyzed with the algorithms ROTCUR and DiskFit, that the gas kinematics can be strongly biased, and a true rotation curve difficult to obtain.
@@ 2 Data
We based our study on a gas-rich giant spiral barred galaxy (gSb) from the GalMer database. This database has been built using smoothed-particle hydrodynamics (SPH) simulations to study the impact of intergalactic interactions on galaxy evolution, but it contains isolated galaxy runs as control samples. For the purpose of our analysis we selected one of these isolated galaxy runs. We chose one with 160 000 particles (run #1030). We wanted well-defined spiral arms and a strong bar so as in Randriamampandry et al. (2016) we analyzed the snapshot at $T=500$ Myrs. From now on, the model will be referred to as gSb. The properties of gSb (mass of the different components, scale lengths and number of gas particles) are given in Table 1.
Table 1: Fundamental parameters for gSb.
@@
$_{\rm{B}}\ [10^{9}MM_{\odot}$] 11.5
$_{\rm{H}}\ [10^{9}MM_{\odot}$] 172.5
$_{*}\ [10^{9}MM_{\odot}$] 46
$_{\rm{g}}\ [10^{9}MM_{\odot}$] 9.2
@@
[…]
Chilingarian et al. (2010)
@@
From the GalMer website [^1] http://galmer.obspm.fr, we obtained masses, positions and velocities of gas particles. We added the same particles at opposite positions (x,y,z)(-x,-y,-z) and opposite velocities $(-v_{x},-v_{y},-v_{z})$ to the galaxy in order to reduce morphological asymmetries, as the different configurations studied are created by rotation, and we want them to differ only by the bar’s orientation. Then, spectral cubes are created only with the gas particles with each pixel measuring 200 pc $\times$ 200 pc and using a spectral resolution of 12 km s-1. The pixel size value represents the minimum softening length used in GalMer (Chilingarian et al., 2010). The bar is about 5 kpc long, which is equivalent to $\sim 25$ pixels. Initially, the galactic disk is in the xyx-y plane and the line-of-sight (LoS) is in the $z$ direction. We create one cube every 10 for an inclination angle $i_{\rm{true}}\in[30^{\circ},80^{\circ}]$, the interval recommended by Di Teodoro & Peek (2021). In order to study symmetrical configurations with respect to the LoS, we vary the inclination orientation which is equivalent to changing the nearest/farthest side. To reproduce observations we smooth the obtained cube using a Gaussian beam with a full width at half maximum (FWHM) of 0.3 kpc (corresponding to $\arcsec=1$ kpc for a galaxy at a distance of $D=206.3$ Mpc). This FWHM corresponds to a maximum baseline of $B_{\rm max}=724$ m at $\nu=350$ GHz, consistent with ALMA CO(3–2) observations in a compact configuration. As an example, this typical emission line was used by Audibert et al. (2019) to observe a trailing spiral inside the ILR ring of the bar in NGC 613.
@@ 4 Kinematics modeling with the bar along the kinematic axes
This allows us to evaluate how these limiting configurations affect 3D-Barolo’s recovery of the geometrical and kinematic parameters plotted in Fig. 2.
Figure 2: Each column corresponds to one bar orientation of Fig. 1. Left column: bar along the minor axis and right column: bar along the major axis. Parameters fitted by 3D-Barolo with a free PA and a free inclination angle. Each marker corresponds to one ’true’ inclination value $i_{\rm{true}}\in\pm[30^{\circ},80^{\circ}]$, the one we input to create the cube. The different parameters are plotted as function of the galactocentric radius and the blue area highlights the bar region. First row: the PA obtained at the end of step 1. Second row: the inclination angles obtained at the end of step 1. Third row: rotation curves obtained after step 2. The black dashed curve represents the expected circular velocity (Eq. 6) computed from the gravitational potential (Eq. 4). Fourth row: Radial velocity with the sign convention: negative towards the center.
@@
We fit both of the geometrical parameters: the PA $\phi$ and the inclination angle $i$ and the kinematic parameters: the tangential velocity $V_{\rm rot}$, the velocity dispersion $V_{\rm disp}$ and the radial velocity $V_{\rm rad}$, plotted in Fig. 2.
@@
Because radial motions project onto the minor axis, the bar–driven radial component is absent in the recovered kinematics when the bar is aligned with the major axis (right column, last row, panel of Fig. 2), whereas a non-zero radial component appears when the bar is along the minor axis (left column, last row, panel of Fig. 2). These amplitudes remain lower than in the $\pm 45^{\circ}$ configurations, explaining our choice of dedicated analysis of those orientations in Sect. 5.
Finally, when radial contribution is almost zero, that is when the bar is along the major axis, 3D-Barolo cannot explain the bar effects by fitting radial velocities, so it shifts the PA by up to 55 within the bar region ($R<2.5\penalty 10000\ \rm kpc)$.
@@ 5 Kinematics modeling with the bar at $\pm 45^{\circ}$ from the kinematic axes
  • a free PA and a fixed inclination angle (Sect 5.3),
  • a fixed PA a fixed inclination angle (Sect. 5.4).
@@ 5.1 Results with a free PA and a free inclination angle
Figure 4: Same as Fig. 2 but each column corresponds to one configuration presented in the rows of Fig. 3.
@@
The lower the inclination, the further the PA estimated by 3D-Barolo in the region $R<4$ kpc is from the expected PA of $\phi=270^{\circ}$ (first row of Fig. 4). This region corresponds to an area larger than the bar region ($R<2.5$ kpc), but errors are significant within this area. Beyond 4 kpc, the computed PA corresponds to the correct value we used as a fixed input in Sect. 5.2, $\phi=270^{\circ}$. However, inside the bar, 3D-Barolo tends to oppose the PA to the bar orientation when it is at $\pm 45^{\circ}$. This is likely due to the influence of the spiral structure around the bar. For a highly inclined disk ($i_{\rm true}=80^{\circ}$ in all configurations and also $i_{\rm true}=70^{\circ}$ in configurations (c) and (d)), 3D-Barolo computes the right PA even in the bar region. The bar is less observable in almost edge-on disks and the algorithm seems to ignore its effect on the geometrical parameters in this case.
@@ 5.3 Results with a free PA and a fixed inclination angle
Fixing the inclination produces a stronger overestimation of the tangential velocity at high inclination in configurations (c) and (d) (second row of Fig. 11). Once the inclination can no longer adjust, the fit compensates for the bar-induced non-circular motions by artificially increasing $V_{\rm rot}$. In the other cases, the rotational curves are similar than the ones obtained with a free $i$ (third row of Fig. 4).
@@
5.4 Results with a fixed PA and a fixed inclination angle
In general, the geometric parameters ($i$, PA) can be determined assuming that the outer regions of the galaxy disks are circular, unperturbed. Therefore, it is interesting to only fit the kinematic parameters with a fixed PA ($\phi=270^{\circ}$) and inclination angle ($i=i_{\rm true}$). The rotation curves and the radial velocities derived by 3D-Barolo with this combination are plotted on Fig. 7.
Figure 7: Same as Fig. 4 but parameters were fitted with a fixed PA of $\phi=270^{\circ}$ and a fixed inclination angle.
The results are very similar to those obtained with a fixed PA and free inclination angle (see Fig. 6), indicating that the PA is the most critical geometrical parameter in the kinematics modeling. We note less discrepancy in the tangential velocities outside the bar region with the varying inclination angle when both geometrical parameters are fixed. One difference from the previous input parameters combinations results is that in configurations (a) and (d) – whose bar’s orientation is the same – two regimes are observed in the rotation curves (first row of Fig. 7) depending on the inclination angle. For an inclination of $30^{\circ}\leq i_{\rm true}\leq 50^{\circ}$, there is a dip feature in the bar region, discussed in Sect. 6.2, whereas for $60^{\circ}\leq i_{\rm true}\leq 80^{\circ}$ it is either non-existent either less pronounced. Similar features appear in the other configurations but with some unexpected values. Even when they are fixed to their true values, we observe a wrong estimation of the rotation curves, particularly within the bar region.
@@ 6 Discussion
The first limitation of our study concerns the algorithm itself. As noted by Di Teodoro & Peek (2021), 3D-Barolo assumes an axisymmetric disk, which is not an adequate description for a barred spiral galaxy. When used on observations, 3D-Barolo usually fits the geometrical parameters of the galactic disk but even in the best cases, where the PA and/or the inclination angle are known and fixed to a constant value, one cannot escape the bar’s effects.
@@
In most cases, when a Bézier curve fits the inclination angle, the mean value computed by 3D-Barolo outside the bar region is closer to the true value when the PA is fixed at $\phi=270^{\circ}$ (first row of Fig. 6) than when it is left free (second row of Fig. 4). However, there are some exceptions in configurations (c) and (d). 3D-Barolo tends to underestimate the median inclination more compared to the values obtained when the PA is free to vary (between 1.9 and 5.1 versus 1.1 and 4.3).
The rotation curves are similar whether the PA is left free (third row of Fig. 4) or fixed (the second row of Fig. 6) outside $R>4$ kpc. At $i_{\rm true}=80^{\circ}$, there are unexpected high values of $V_{\rm rot}$ in configurations (b) and (d) (second row of Fig. 6) around $R\sim 3$ kpc that are not there when the PA is free (third row of Fig. 4) as the computed inclination was a constant median value.
The apparent radial velocities obtained by 3D-Barolo in the bar region when the PA is left free (fourth row of Fig. 4), are lower than with a fixed PA of $\phi=270^{\circ}$ (third row of Fig. 6).
@@
For highly inclined disk, the apparent radial velocities decrease, reaching values close to 0 for $i_{\rm{true}}=80^{\circ}$ in certain cases. The values change a lot with the configurations and the input parameters combination. This confirms that the bar affects the recovery of both geometrical and kinematic parameters, introducing substantial uncertainty in the apparent radial velocities inferred by 3D-Barolo. The low radial velocities force the model to absorb the bar-driven non-circular motions into the tangential component, producing the artificially high values of $V_{\rm{rot}}$ at this inclination (second rows of Fig. 6, Fig. 11 and first row of Fig. 7). At this high inclination, 3D-Barolo attempts to fit circular orbits, effectively ignoring the non-circular motions produced by the bar.